OpenStudy (anonymous):

A 1.2 x 10^3 kg car accelerates uniformly from 5.0 m/s east to 12 m/s east. During this acceleration the car travels 94 m. What is the net force acting on the car during the acceleration? I know F=mass x acceleartion. We have the mass so I need to find the acceleration first. I know acceleration is velocity over time. So 12 m/s - 5 m/s = 7 m/s. We need the time, and I know that time is distance over velocity, so 94m / 7m/s = 13s. Now I can try to find the acceleration. 7m/s / 13s = 0.54 m/s^2. So force = 1200 kg x 0.54 m/s^2. I get 648 N, but the answer is 760 N. What am I doing wrong

OpenStudy (danjs):

do you remember what the kinematic equations are

OpenStudy (danjs):

there is a trick in figuring those out, where the time variable falls out of one of those equations, in terms of other variables

OpenStudy (danjs):

when you are deriving them

OpenStudy (danjs):

you remember this looking thing? combo of the others \[Vf^2=Vo^2+2a*(Xf-Xo)\]

OpenStudy (lochana):

" so 94m / 7m/s = 13s" <-- this is not correct

OpenStudy (lochana):

@DanJS please continue:)

OpenStudy (danjs):

you can solve directly from that.. given the change in velocity, and the change in distance, you can figure the acceleration... (constant)

OpenStudy (danjs):

Accelerated from 5 to 12 m/s over a distance of 94m

OpenStudy (danjs):

remember, the acceleration is a CHANGE in a velocity over a time , you put 7 m/s, the change in the velocity , in to the distance formula d=v*t, in the distance formula, that is a constant velocity value, not the change in 2 values

OpenStudy (danjs):

oh, i was waiting for you to respond @lochana , i just now realized this is not your original question.. ... hahah

OpenStudy (lochana):

okay. so we can use \[V^2 = U^2 + 2as \] for this problem right? http://www.schoolphysics.co.uk/age14-16/Mechanics/Motion/text/Equations_of_motion/index.html

OpenStudy (lochana):

u = initial velocity which is 7m/s v = last velocity which is 12m/s s = distance that is 94 a = acceleration

OpenStudy (lochana):

yes. that's right

OpenStudy (lochana):

so \[a = \frac{V^2 - V0^2}{2(x-x0)}\]\[a = \frac{144 - 25}{2*94}\]

OpenStudy (danjs):

a constant accel , means you can write the acceleration definition a=change in V / Change in Time as the first kinematic equation a*t = Vf - Vi

OpenStudy (danjs):

since a is taken as a constant

OpenStudy (lochana):

yes. but our goal is to find 'a'. we still don't know about 't' (time).

OpenStudy (lochana):

\[a = \frac{144 -25}{2*94} = \frac{119}{188}ms^{-2}\] this is the constant acceleration. m is given (1.2 x 10^3 kg)

OpenStudy (danjs):

if acceleration is constant, the velocity changes at a uniform rate... remember <V> the average velocity.. <V>=change in velocity over 2

OpenStudy (lochana):

\[F_{net} = ma\]

OpenStudy (lochana):

@DanJS do you think my solution for 'a' is correct?

OpenStudy (danjs):

THe average velocity is also the total distance over the total time <V> = (Xf - Xi)/(tf - ti)

OpenStudy (lochana):

can you show steps to find "a"?

OpenStudy (danjs):

combine those two , and you get the next kinematic thing to remember, \[<V >= \frac{ \Delta x }{ \Delta t } = \frac{ \Delta V }{ 2 }\]

OpenStudy (danjs):

i found what I was typing.. Follow this, see how you can figure out that kinematic equation relating Velocity, Distance, and Acceleration

OpenStudy (danjs):

There are other ways, this is straight forward from other ideas of accel and velocity and position http://theory.uwinnipeg.ca/physics/onedim/node7.html#eq:_first_EOM

OpenStudy (lochana):

@Openstudystudent314 I didn't notice that you have given the solution to your question. look at my answer. you get a = 119/188. now multiply that with 1200kg. you get 760N.

OpenStudy (lochana):

@DanJS yes. what you are saying is correct. we can solve this question in many ways.

OpenStudy (danjs):

i was trying to show how you can derive that equation i gave, in the process it shows the time variable dropping out, and you can relate accel to distance and change in velocity

OpenStudy (danjs):

constant accel, uniform change in velocity

OpenStudy (lochana):

yeah. I remember that. it takes some extra steps to solve it. otherwise both are same:)

OpenStudy (danjs):

i think we used the Force = dp/dt to start, not sure.. long time back

OpenStudy (danjs):

change in momentum